OpenStudy (anonymous):

The ordered pair (-1, -1) is a solution of the system -4x + 2y = 2 and x + y = -2. True False

OpenStudy (anonymous):

true

OpenStudy (anonymous):

true .....points stasify

OpenStudy (anonymous):

Eq one is -4x + 2y = 2 Eq two is x + y = -2. Now solve two eq's  Putting the value of y in eq no. one we get -4x+2(-1)=2 -4x-2=2 -4x=2+2 -4x=4 x=-1 So the solution of the equation is (-1,-1) and answer is true